# TRA: Better Length Generalisation with Threshold Relative Attention

### v3

@@ 4 Related Work:
Generalisation in Neural Networks: Neural networks, and particularly transformers, struggle with out of distribution generalisation (Hupkes et al., 2019; 2023). Partially this is architecture dependent. Significant efforts have been made towards creating architectures that enable a more flexible information flow and consequently have access to greater space of possible solutions (Fan et al., 2020; Chowdhury & Caragea, 2021; Opper et al., 2023; Soulos et al., 2024; DuSell & Chiang, 2024). However, these approaches tend to come at the cost of either flexibility or parallelisability - the key properties which made transformers, and particularly their decoder only variants, take over the world.
@@
Transformer Programs: A parallel effort has been made to understand what kind of finite depth solutions transformers should be able to represent (Weiss et al., 2021; Friedman et al., 2023; Smolensky et al., 2024). One motivation being the conjecture that transformers will learn to generalise iff they are able to represent a program which does so (Zhou et al., 2023). However, empirical validation for this conjecture has thus far remained limited. Even when we expect the models to generalise they often only do so weakly, and rigorously bridging the gap between theory and practice remains a long standing challenge. A key motivation for TRA came from the theoretical model presented by Smolensky et al. (2024) which draws analogy between the transformer and classical production system architecture from cognitive science. Their theoretical model has the attention mechanism first exactly match on specified conditions before applying a hard positional tie-breaker to perform one-hot selection. TRA arose out of an ongoing effort to create a learnable equivalent to their theoretical model.
@@
Length Generalisation in Transformers: Several mechanisms have been posisted to improve transformer length generalisation. We partition this work into two categories. The first deals with trying to learn positional weights that are flexible w.r.t. the length distribution of the training data (Dubois et al., 2019; Ruoss et al., 2023; Li & McClelland, 2023; Ray Chowdhury & Caragea, 2023; Golovneva et al., 2024). Orthogonally, recent work has demonstrated that when extending to longer sequences the softmax tends towards uniformity and therefore introduces errors the attention signal (Ye et al., 2024; Veličković et al., 2024). TRA is intended to demonstrate that these two streams of research should in fact combine. It both attempts to mitigate noise in the attention through selectively sparsity and also learn general purpose positional information through its forget gate.
@@ 5 Conclusion
Appendix A Hyperparameters:
For RoPE we set the theta to 500k following Dubey et al. (2024). For CoPE we set npos max to 64 following the original authors Golovneva et al. (2024). All models are trained using the AdamW optimiser with a low dropout of 0.01 applied to the attention weights and feedforward hidden representation.
Appendix B Implementation:
1class TRA(nn.Module):
@@ Appendix B Implementation:
7 self.tokeys = nn.Linear(e_dim, e_dim, bias=False)
8 self.toqueries = nn.Linear(e_dim, e_dim, bias=False)
9 self.tovalues = nn.Linear(e_dim, e_dim, bias=False)
10 self.delta_proj = nn.Linear(e_dim, self.h)
11 self.c_out = nn.Linear(e_dim, e_dim)
@@
15 self.dropout = nn.Dropout(dropout)

### v4

@@ 4 Related Work:
Generalisation in Neural Networks: Neural networks, and particularly transformers, struggle with out of distribution generalisation (Hupkes et al., 2019; 2023). Partially this is architecture dependent. Significant efforts have been made towards creating architectures that enable a more flexible information flow and consequently have access to greater space of possible solutions (Fan et al., 2020; Chowdhury & Caragea, 2021; Opper et al., 2023b; Opper & Siddharth, 2024b; Soulos et al., 2024; DuSell & Chiang, 2024). However, these approaches tend to come at the cost of either flexibility or parallelisability - the key properties which made transformers, and particularly their decoder only variants, take over the world.
@@
Transformer Programs: A parallel effort has been made to understand what kind of finite depth solutions transformers should be able to represent (Weiss et al., 2021; Friedman et al., 2023; Smolensky et al., 2024). One motivation being the conjecture that transformers will learn to generalise iff they are able to represent a program which does so (Zhou et al., 2023). However, empirical validation for this conjecture has thus far remained limited. Even when we expect the models to generalise they often only do so weakly, and rigorously bridging the gap between theory and practice remains a long standing challenge. A key motivation for TRA came from the theoretical model presented by Smolensky et al. (2024) which draws analogy between the transformer and the classical production system architecture from cognitive science. Their theoretical model has the attention mechanism first exactly match on specified conditions before applying a hard positional tie-breaker to perform one-hot selection. TRA arose out of an ongoing effort to create a learnable equivalent to their theoretical model.
@@
Length Generalisation in Transformers: Several mechanisms have been posisted to improve transformer length generalisation. We partition this work into two categories. The first deals with trying to learn positional weights that are flexible w.r.t. the length distribution of the training data (Dubois et al., 2019; Ruoss et al., 2023; Li & McClelland, 2023; Ray Chowdhury & Caragea, 2023; Golovneva et al., 2024). Orthogonally, recent work has demonstrated that when extending to longer sequences the softmax tends towards uniformity and therefore introduces errors the attention signal (Ye et al., 2024; Veličković et al., 2024). TRA investigates whether these two streams of research should combine. It attempts to reduce noise in the attention through selectively sparsity and learns general positional encodings with its forget gate.
Sparsifing Attention: Finally, a substantial amount of effort has been made towards introducing sparsity into attention. One vein attempts to reduce the quadratic cost of the attention operation by applying fixed patterns (Child et al., 2019; Zaheer et al., 2021; Nawrot et al., 2025), though doing so either requires restricting flexibility or is only applicable during inference time. Another has sought to allow sparsity within the softmax by allowing for zero probabilities (Peters et al., 2019), however, such algorithms often depend on computational complex operations and have delivered mixed results. Finally, Zhang et al. (2021) replace the softmax with ReLU, thereby introducing selective sparsity and demonstrating promising results on NMT.
@@ 5 Conclusion
Appendix A Implementation:
1class TRA($n$.Module):
@@ Appendix A Implementation:
7 self.tokeys = $n$.Linear(e_dim, e_dim, bias=False)
8 self.toqueries = $n$.Linear(e_dim, e_dim, bias=False)
9 self.tovalues = $n$.Linear(e_dim, e_dim, bias=False)
10 self.delta_proj = $n$.Linear(e_dim, self.h)
11 self.c_out = $n$.Linear(e_dim, e_dim)
@@
15 self.dropout = $n$.Dropout(dropout)
@@ 37 A[mask] = -1e11 # mask: casual + threshold
Appendix B Hyperparameters:
For RoPE we set the theta to 500k following Dubey et al. (2024). For CoPE we set npos max to 64 following the original authors (Golovneva et al., 2024). All models are trained using the AdamW optimiser with a low dropout of 0.01 applied to the attention weights and feedforward hidden representation.
Appendix C Attention Analysis: Language Modeling
Trained TRA models exhibit both completely content focused and completely positional attention heads, as well as ones where the trade-off is somewhere in between. As a general rule of thumb the more sparsely activated the QK content weights, the weaker the degree of time decay enforced by the forget gate. Conversely, dense QK activations lead the model to focus on a local window (see Figure 3 below). Attention patterns can be highly sparse indicating significant threshold pruning. Secondly, tokens can exhibit considerable variance in time decay, impossible with fixed schemes such as ALiBi Press et al. (2021).
(a) Densely Activated Attention Head: Position Outweighs Content.
(b) Mixed Activation Head: Content and Position Balanced.
(c) Sparsely Activated Attention Head: Content Dominates Position
Figure 3: Attention Heatmaps from trained TRA language model. White signifies null attention.
Appendix D Computational Efficiency and Potential Optimisations:
One potential disadvantage of TRA compared with schemes such as RoPE, which remain the current defacto choice for positional embeddings, is that it introduces additional operations which can diminish runtime efficiency. To give an indication across two configurations:
Tiny: 4 layers, 4 heads, residual stream size 256, ffn 1024
Medium: 8 layers, 8 heads, residual stream size 512, ffn 1024
On a single Nvidia A40 card with window size 256 and batch size 64:
Tiny RoPE: 20k steps 58.73 minutes
Tiny TRA: 20k steps 65.88 minutes (circa seven minute increase)
Medium RoPE: 20k steps 156.34 minutes
Medium TRA: 20k steps 178.19 minutes (circa 22 minute increase)
However, these differences are without any effort towards runtime optimisation on our end. We believe there is significant scope for even greater efficiency. Though this is currently highly speculative, and hence left out of the main paper, there is no reason why the TRA algorithm could not be computed in a tiled manner and take advantage of the optimisations introduced by flash attention Dao et al. (2022); Dao (2023). While not trivial because the blocks we need to be processed in a causal manner to allow for online computation of the cumsum, the fact that attention patterns can be highly sparse and the forget gate can enable early stopping of row processing means that there should be considerable scope for greatly improving the efficiency of TRA.